OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope =5/3 ; y-intercept = (0, 2) slope = -3/5; y-intercept = (0, 2) slope = 5/3; y-intercept = (0, -6)

OpenStudy (anonymous):

We have to isolate y so we subtract 5x on both sides. -3y=-5x-6 and now we divide both sides by -3. y=5/3x+2 and the answer is B.

OpenStudy (anonymous):

sorry it's b.

OpenStudy (anonymous):

\[5x-3y=-6\rightarrow y=(-6-5x)/(-3)\rightarrow y=5x/3+2\]Thus slope is 5/3 and y-inter is 2

OpenStudy (anonymous):

m=rise over run